### v1

# GALD-SE: Guided Anisotropic Lightweight Diffusion for Efficient Speech Enhancement

Abstract
Speech enhancement is designed to enhance the intelligibility and quality of speech across diverse noise conditions. Recently, diffusion model has gained lots of attention in speech enhancement area, achieving competitive results. Current diffusion-based methods blur the initial signal with isotropic Gaussian noise and recover clean speech from the prior. However, these methods often suffer from a substantial computational burden. We argue that the inefficiency stems from the oversight that speech enhancement is not purely a generative task; it primarily involves noise reduction and completion of missing information, while the clean clues in the original mixture do not need to be regenerated. In this paper, we propose a method that introduces noise with anisotropic guidance during the diffusion process, allowing the neural network to focus on clean clues within noisy recordings. This approach is robust against various types of noise interference and speech distortion, and significantly reduces the computational load. Experiments demonstrate that the proposed method achieves state-of-the-art results with only approximately 4.5 million parameters, significantly fewer than those required by other diffusion methods. This effectively narrows the gap in model size between diffusion-based and predictive speech enhancement approaches. Additionally, the proposed method performs well in very noisy scenarios, demonstrating its potential for applications in highly challenging environments.
Index Terms:
Speech enhancement, generative diffusion models, anisotropic diffusion process
I Introduction
Speech enhancement aims at recovering clean speech signal from a noisy mixture that is corrupted by various types of noise. It has a wide range of applications, including hearing aids, speech recognition systems, and human-computer interaction. Traditional methods usually involve noise power estimation and extracting the clean signal in a statistical signal processing way [1, 2, 3]. Not long ago, numerous machine learning-based speech enhancement techniques were proposed [4, 5, 6, 7]. These methods involve training a neural network to predict the clean signal from its noisy counterpart, significantly surpassing traditional methods in terms of speech quality, especially in non-stationary environments [8]. Recently, diffusion-based generative models have gained increasing attention [9][23], achieving state-of-the-art results and surpassing predictive approaches when the tested audio mismatches with training datasets [9], [16].
Diffusion-based methods operate through a two-phase process: gradually adding noise to an initial state in the forward process, and then iteratively sampling along the reverse process to generate targeted samples. In the field of speech enhancement, the spectrogram of noisy speech frequently serves as a component of the prior for the diffusion process and as an input to the neural network for reverse sampling [9], [10], [11], [13], or exclusively the latter [17, 18, 22]. The corresponding diffusion process can be interpreted in a discrete [29] or continuous [9] way, while the latter permits the use of numerical stochastic differential equation (SDE) solvers to generate estimated speech [25]. Richter et al. [9] proposed an SDE-based method that introduces a novel drift term. In this way, the mean state of the process is transformed from clean speech into its noisy counterpart during the forward process and reversed during sampling. Two main approaches have been explored to improve upon this scheme: (i) Carefully designing new interpolation techniques, such as the Brownian bridge [15, 16] and variance-preserving interpolation methods [13, 14], which have demonstrated significant performance improvements; and (ii) Integrating diffusion models with predictive methods [12, 20, 23]. The latter approach, incorporating a predictive model, can sometimes reduce the computational complexity of the original diffusion method. However, even the smallest version requires approximately 27 million parameters to achieve satisfactory results [23], making diffusion-based methods highly computationally intensive. We argue that this burden stems from the inefficiency of the diffusion process. Current diffusion scheme doesn’t effectively make use of the clean clues in the mixture and isotropically blur them during the forward process. This blurring distorts the original clean clues, forcing the neural network to re-estimate them during the reverse process, thereby increasing the computational load.
Based on these insights, we propose GALD-SE, a Guided Anisotropic Lightweight Diffusion-based Speech Enhancement method, which incorporates anisotropic guidance in the diffusion process. The guidance effectively preserves the integrity of clean clues within noisy mixtures, enabling the process to directly access these clues at any timestep. The formulation of anisotropic diffusion process is built upon the foundational work presented in [27, 28]. Experimental results demonstrate that our diffusion-based method achieves state-of-the-art performance while utilizing only 4.5 million parameters, which is less than 20% of the parameter count in the smallest existing diffusion model [23]. Ablation analysis and comparison confirm the efficiency and superior quality of generated speech across various performance metrics.
II Proposed Method
The system overview of our method is depicted in Fig. 2. An anisotropic noise guidance is derived from a coarse magnitude estimation network. This guidance directs the diffusion process to blur each time-frequency (T-F) bin with specific intensity. During each step of the reverse process, the diffusion network takes the current state, anisotropic guidance, and the noisy speech as inputs. It then computes an estimate of the initial state, which corresponds to the clean spectrogram. This estimated initial state is then used to derive the state at the previous timestep, generating the final result through multiple iterations. By implementing anisotropic guidance to minimize unnecessary diffusion, we enhance both the efficiency and performance. The rest of this section is organized as follows: First, we explain the modifications made to the original diffusion process. Next, we provide the calculation of anisotropic guidance. Finally, we describe the training and sampling strategies of our method.
II-A Stochastic Process
We denote the complex STFT feature vectors of clean and noisy speech as $\mathbf{x}_{0}$ and $\mathbf{y}$, respectively. These vectors reside in $\mathbb{C}^{d}$, where $d=KF$, representing the coefficients of a flattened complex spectrogram. In the typical “ResShift” formulation [27], during the forward process, the mean of the state gradually transitions from $\mathbf{x}_{0}$ to $\mathbf{y}$ by adding isotropic noise at each step. The conditional distribution of the state at time $t$, given the state at time $t-1$, the initial state $\mathbf{x}_{0}$, and the noisy feature $\mathbf{y}$, is given by:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{x}_{0},\mathbf{y})=\mathcal{N}_{\mathbb{C}}\left(\mathbf{x}_{t};\mathbf{x}_{t-1}+\alpha_{t}(\mathbf{y}-\mathbf{x}_{0}),\kappa^{2}\alpha_{t}\mathbf{I}\right),$ (1)
$t=1,...,T$
where $\mathcal{N}_{\mathbb{C}}$ denotes the circularly-symmetric complex normal distribution, $\mathbf{x}_{t}$ represents the state at timestep $t$, $\kappa$ is a hyperparameter controlling the noise variance, $\mathbf{I}\in\mathbb{R}^{KF\times KF}$ is the identity matrix, and $\alpha_{t}$ reflects the noise schedule. We contend that the isotropic Gaussian noise added at each step can gradually impair the clean clues in the signal. Therefore, we propose to update each step of the forward process as follows:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{x}_{0},\mathbf{y})=\mathcal{N}_{\mathbb{C}}\left(\mathbf{x}_{t},\mathbf{x}_{t-1}+\alpha_{t}(\mathbf{y}-\mathbf{x}_{0}),\kappa^{2}\alpha_{t}\boldsymbol{\sigma}^{2}\right)$ (2)
$t=1,...,T$
where $\boldsymbol{\sigma}\in\mathbb{R}^{KF\times KF}$ is the anisotropic noise guidance matrix, designed to preserve the clean clues within noisy mixture. This matrix is diagonal, resulting in Gaussian noise that is independent across different T-F bins. During the forward process, the real and imaginary components of the complex features within each T-F bin undergo independent diffusion with the same variance, which is decided by the estimated proportion of noise in that bin. An illustrative example of the anisotropic process is depicted in Fig. 2(a). Suppose that in the original mixture, the 2nd, 3rd, and 5th frequency bins contain a smaller proportion of noise and should therefore be blurred less, while the 1st, 4th, and 6th bins contain a larger proportion of noise and should undergo more blurring. Consequently, for the 2nd, 3rd and 5th bins, the Gaussian noise added at each step is relatively small, keeping the values at these spectral positions nearly unchanged. In contrast, for the other bins, the Gaussian noise added at each timestep is much larger. In this way, the diffusion process can selectively blur T-F bins with varying intensities during the forward process, thereby preventing the introduction of excessive noise to bins that predominantly contain clean speech components.
Although the anisotropic guidance introduces varying intensities of noise to different bins, the overall mean of the process remains unaffected due to the zero-mean nature of the Gaussian noise. Consequently, following the derivation in [27], we can directly sample $\mathbf{x}_{t}$ from the initial state $\mathbf{x}_{0}$ as follows:
$q\left(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{y}\right)=\mathcal{N}_{\mathbb{C}}\left(\mathbf{x}_{t};(1-\overline{\alpha}_{t})\mathbf{x}_{0}+\overline{\alpha}_{t}\mathbf{y},\kappa^{2}\overline{\alpha}_{t}\boldsymbol{\sigma}^{2}\right),$ (3)
$t=1,...,T$
where $\overline{\alpha}_{t}=\Sigma_{k=1}^{t}\alpha_{k}$. The corresponding reverse process can be depicted as:
$q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{y}\right)=\mathcal{N}_{\mathbb{C}}\left(\mathbf{x}_{t-1}|(1-\beta_{t})\mathbf{x}_{t}+\beta_{t}\mathbf{x}_{0},\overline{\boldsymbol{\sigma}}_{t}^{2}\right)$ (4)
where $\beta_{t}=\frac{\alpha_{t}}{\overline{\alpha}_{t}}$ and $\overline{\boldsymbol{\sigma}}_{t}^{2}=\kappa^{2}\beta_{t}(1-\beta_{t}){\boldsymbol{\sigma}}^{2}$. The calculation of the anisotropic guidance $\boldsymbol{\sigma}$ will be detailed in the following section.
II-B Anisotropic Guidance
We developed anisotropic guidance specifically to blur the noise structure while minimally altering the clean components. To achieve this, we consider the ideal ratio mask (IRM), which represents the proportion of magnitude of clean components present in each T-F bin. By utilizing a lightweight UNet model, denoted as $\mathbf{g}_{\delta}(\mathbf{y})$, which is adapted from [30] and parameterized by $\delta$, we input the noisy feature $\mathbf{y}$ to obtain a clean magnitude estimation. Subsequently, we derive a coarse IRM, denoted as $\tilde{\mathbf{M}}_{0}$, residing in $\mathbb{R}^{K\times F}$:
$\tilde{\mathbf{M}}_{0}=\mathbf{g}_{\delta}(\mathbf{y})/\left\|\mathbf{y}\right\|_{2}$ (5)
We then define the guidance matrix $\boldsymbol{\sigma}$ to be diagonal, with its elements obtained as follows:
$\boldsymbol{\sigma}=\operatorname{diag}(\operatorname{vec}(\mathbf{1}-\tilde{\mathbf{M}}_{0}))$ (6)
where $\mathbf{1}$ is a matrix in $\mathbb{R}^{K\times F}$ with all elements set to 1. It is evident that the guidance $\boldsymbol{\sigma}$ is derived from the proportion of noise present in each T-F bin. The higher the noise ratio in a bin, the greater the corresponding value of $\sigma$ is in that bin. In this way, we encourage that the noise in the origin mixture loses its original structure during the forward process, while the clean structure remains nearly intact. This guidance simplifies the model’s task of distinguishing clean speech from the mixture and reduces the risk of receiving misleading clues from the noise parts.
Figure 1: System overview.
Figure 2: Visualization of anisotropic diffusion process and its impact on speech enhancement. (a) Illustrative example of the anisotropic diffusion process. (b) Impact of anisotropic guidance on the prior state and final enhanced output.
We further give an visualization for the impact of anisotropic noise guidance as in Fig. 2(b). The top row displays the noisy spectrogram, followed by the prior sampling states without and with guidance, respectively. The second row indicates the corresponding clean speech and the enhanced results. We observe that with anisotropic guidance, the T-F structure of the noise is significantly blurred, while the initial state more effectively preserves clean speech clues compared to its counterpart without guidance. This preservation facilitates the model’s estimation of the original clean spectrogram. Consequently, the enhancement result with anisotropic guidance retains more speech details than those without guidance.
II-C Training and Sampling
At training stage, we sample a state at an arbitrary time $t$ as described in (3), incorporating guidance from (6). The diffusion network, denoted as $\mathbf{f}_{\theta}(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t)$, is parameterized by $\theta$ and processes inputs including the current state $\mathbf{x}_{t}$, noisy feature $\mathbf{y}$, anisotropic guidance $\boldsymbol{\sigma}$, and the current timestep $t$. We design the model to estimate the initial state $\mathbf{x}_{0}$, as the presence of anisotropic guidance complicates the direct estimation of Gaussian noise. This model, along with the coarse magnitude estimation network $\mathbf{g}_{\delta}(\mathbf{y})$ discussed in Section II-B, are simultaneously trained using a mean square error criterion, leading to the following optimization objective:
$\min_{\boldsymbol{\theta},\boldsymbol{\delta}}\sum_{t}\left[\left\|\mathbf{f}_{\theta}\left(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t\right)-\mathbf{x}_{0}\right\|_{2}^{2}+\left\|\mathbf{g}_{\delta}(\mathbf{y})-\left\|\mathbf{x}_{0}\right\|_{2}\right\|_{2}^{2}\right]$ (7)
At sampling stage, we first employ (5), (6) to derive the anisotropic guidance $\boldsymbol{\sigma}$, and then get $\overline{\boldsymbol{\sigma}}_{T}$ based on (4). To initiate the reverse sampling procedure, we sample the initial state as follows:
$\mathbf{x}_{T}=\mathbf{y}+\overline{\boldsymbol{\sigma}}_{T}\mathbf{z},\quad\mathbf{z}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{z};0,\mathbf{I})$ (8)
At each reverse step $t$, we derive the previous state $t-1$ as follows:
$\mathbf{x}_{t-1}=(1-\beta_{t})\mathbf{x}_{t}+\beta_{t}\tilde{\mathbf{x}}_{0}+\overline{\boldsymbol{\sigma}}_{t}\mathbf{z},\quad\mathbf{z}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{z};0,\mathbf{I})$ (9)
where $\beta_{t}$ and $\overline{\boldsymbol{\sigma}}_{t}$ retain their definitions as in (4). By applying this reverse sampling process, we iteratively derive the final estimation $\mathbf{\tilde{x}}_{0}$ from the initial state $\mathbf{x}_{T}$.
III Experiments
III-A Datasets and Evaluation Metrics
We use DNS-Challenge [31] and VoiceBank+Demand (VBD) [32] as the verified datasets. For the DNS corpus, we generated 1,000,000 pairs for training and 300 pairs for validation, each lasting 6 seconds, with signal-to-noise ratios (SNR) ranging from -15 to 25 dB. To construct the test set, we randomly selected 1000 English read speech clips, distinct from those in the training set, and 1000 noise clips, each with a duration of 6 seconds. These clips were mixed at various SNR levels ranging from -15 dB to 15 dB, divided into three intervals. The VBD corpus, widely used as a benchmark dataset in speech enhancement, consists of a training set with 28 speakers subjected to 8 noise types across 5 SNR levels, and a test set featuring 2 speakers with unseen noise types.
We choose the following widely used metrics to evaluate the performance: 1) PESQ: the Perceptional Evaluation of Speech Quality [33]; 2) ESTOI: the Extended Short-Time Objective Intelligibility [34]; 3) SI-SNR: Scale-Invariant Signal-to-Noise Ratio; 4) CSIG, CBAK, COVL: the speech quality (CSIG), background noise quality (CBAK) and the overall quality (COVL) of the audio [35] which reflects the mean opinion scores; 5) DNSMOS: A non-intrusive perceptual objective speech quality metric to evaluate noise suppressors [36].
TABLE I: Performance comparison of different methods on DNS Dataset
Metrics Type Para. Steps PESQ ESTOI SI-SNR (dB) DNSMOS
SNR(dB) -15$\sim$-5 -5$\sim$5 5$\sim$15 Avg. -15$\sim$-5 -5$\sim$5 5$\sim$15 Avg. -15$\sim$-5 -5$\sim$5 5$\sim$15 Avg. -15$\sim$-5 -5$\sim$5 5$\sim$15 Avg.
Noisy - - - 1.17 1.17 1.44 1.26 0.37 0.56 0.75 0.56 -9.96 -0.36 9.65 -0.22 2.46 2.68 3.00 2.71
DEMUCS [37] D 33.5M - 1.27 1.74 2.46 1.82 0.48 0.76 0.87 0.70 2.33 12.28 18.23 10.94 2.88 3.32 3.60 3.27
CMGAN [38] D 1.83M - 1.54 2.11 2.77 2.14 0.62 0.80 0.88 0.77 3.11 7.33 13.75 8.06 3.04 3.43 3.66 3.38
DeepFilterNet [39] D 1.80M - 1.32 1.84 2.64 1.93 0.50 0.74 0.85 0.70 3.28 10.50 16.43 10.07 2.79 3.22 3.61 3.21
SGMSE+ [9] G 65M 30*2 1.38 1.99 2.78 2.05 0.53 0.80 0.88 0.74 -1.43 11.10 18.12 9.27 3.57 3.87 3.92 3.79
StoRM [12] D+G 55.6M 15*2 1.43 1.96 2.73 2.04 0.59 0.80 0.88 0.76 3.97 12.23 18.63 11.61 3.80 3.92 3.94 3.89
VPIDM [13] G 65M 25*2 1.55 2.13 2.88 2.19 0.62 0.83 0.88 0.78 6.17 13.18 19.08 12.81 3.62 3.88 3.95 3.82
BBED [16] G 65M 30*2 1.58 2.14 2.91 2.21 0.65 0.83 0.89 0.79 7.34 14.13 19.94 13.81 3.58 3.84 3.92 3.78
GALD-SE (Ours) D+G 4.5M 10 1.57 2.20 2.96 2.25 0.62 0.81 0.89 0.77 6.44 13.07 18.71 12.74 3.73 3.92 3.94 3.86
III-B Hyperparameters and Training Settings
For the stochastic process, we set the number of diffusion steps $T=10$ and the noise variance strength $k=0.19$ in (2). The noise schedule $\overline{\alpha}_{t}$ is bounded between $\overline{\alpha}_{1}=0.001$ and $\overline{\alpha}_{T}=0.999$. A geometric schedule is adopted for intermediate steps: $\sqrt{\overline{\alpha}_{t}}=\sqrt{\overline{\alpha}_{1}}\times\left(\sqrt{\frac{\overline{\alpha}_{T}}{\overline{\alpha}_{1}}}\right)^{\frac{t-1}{T-1}}$. We adapted the diffusion network architecture from SGMSE+ [9]. The feature dimension is set to 32, resulting in 3.6 million parameters for the diffusion model. The coarse magnitude estimation network, adapted from [30], uses only a magnitude mask decoder, contributing 0.9 million parameters. Consequently, the entire system comprises 4.5 million parameters.
For the STFT configuration, we employ a 510-point FFT with a hop size of 128 samples. Following [9], we compress the resulting spectrum using an exponential factor of 0.5 and then scale it by a factor of 0.15. For training, we set the batch size to 15 and use a learning rate of $10^{-4}$. The model is trained for 500 epochs on the VBD corpus and 30 epochs on the DNS corpus.
III-C Results
We initially compare our method with others using the DNS corpus, as shown in Table I, where ‘D’ denotes discriminative methods and ‘G’ represents generative approaches. Our results demonstrate that the proposed method achieves performance comparable to state-of-the-art techniques across multiple evaluation metrics. Although our method shows a slight performance gap in ESTOI and SI-SNR metrics compared to the best-performing BBED approach, it achieves this with significantly fewer parameters (4.5M vs. 65M) and sampling steps (10 vs. 60). We attribute this significant computational efficiency to our novel implementation of anisotropic noise guidance, which preserves clean speech clues and streamlines the diffusion process. It is worth noting that PESQ may lack reliability in low-SNR scenarios, where ESTOI emerges as a more suitable metric. Our results demonstrate that performance improvements are substantially more pronounced in low-SNR conditions compared to high-SNR situations (an ESTOI improvement of 0.25 in the -15 to -5 dB range versus 0.14 in the 5 to 15 dB range). This underscores the efficacy of our method in challenging acoustic environments.
TABLE II: Performance Comparison of Different Methods on VoiceBank/DEMAND Dataset
Methods Type PESQ ESTOI CSIG CBAK COVL
Mixture - 1.97 0.79 3.35 2.44 2.63
NCSN++ [24] D 2.73 0.86 3.45 3.39 3.17
UNet [30] D 2.87 0.87 4.20 3.48 3.60
DEMUCS [37] D 3.07 - 4.31 3.40 3.63
CDiffuse [11] G 2.52 - 3.72 2.91 3.10
SGMSE+ [9] G 2.93 0.87 4.12 3.37 3.51
StoRM [12] D+G 2.93 0.88 - - -
Thunder [23] D+G 2.97 0.87 - - -
VPIDM [13] G 3.16 0.87 4.23 3.53 3.70
GALD-SE (Ours) D+G 3.19 0.88 4.25 3.59 3.73
   –w/o guidance G 2.93 0.85 4.10 3.15 3.48
   –w/o adding noise G 2.79 0.83 3.97 2.99 3.35
Subsequently, we evaluate our method against other state-of-the-art approaches on the VBD dataset. The results are presented in Table II. The NCSN++ approach adapts the diffusion model architecture to function as a discriminative backbone, whereas the UNet approach exclusively employs the first-stage coarse magnitude estimation, leaving the phase unaltered. Our results demonstrate that the combination of diffusion with anisotropic guidance outperforms approaches using either the coarse magnitude frontend or diffusion as a discriminative backbone alone. Moreover, when compared to other state-of-the-art methods, our proposed approach achieves superior performance across most of the metrics. We posit that this advantage stems from our novel strategy of deeply integrating clean speech clues into the diffusion process through guidance, which enables a more nuanced and effective speech enhancement process.
Finally, we conducted an ablation study to demonstrate the effectiveness of our anisotropic guidance method, as shown in the last three rows of Table II. We compared our proposed method against two variants: One without anisotropic guidance and another without adding Gaussian noise entirely. The latter case, where Gaussian noise is absent throughout the process, corresponds to the cold diffusion formulation [26]. Results indicate that incorporating an appropriate amount of Gaussian noise during the process, rather than directly corrupting clean speech with original noise, yields better results. Moreover, applying guidance to the Gaussian noise during diffusion significantly enhances performance across all evaluated metrics, thus validating the effectiveness of our proposed method.
IV Conclusion
This letter introduces a novel diffusion-based speech enhancement method that leverages guided anisotropic noise. By focusing on the preservation of clean speech clues in noisy environments, our approach reduces the computational demands typically associated with diffusion-based methods, and achieves superior performance. Further reduction of computational load and exploration of the potential for real-time audio stream processing are currently under investigation.